# Assessing (im)balance in signed brain networks

### v1

@@ VIII.3.1 A survey on individual projections
Let us start by considering a pair of individual partitions, stressing that a large link density does not prevent BIC minimisation from detecting statistically significant mesoscopic structures: its sensitivity to both the number and the sign of links, in fact, makes it capable of partitioning both denser and sparser configurations.
Naïve projections. As figs. 10 and 11 show, BIC minimisation reveals the presence of less than 101010 (more precisely, from 555 to 666) communities: the vast majority of links within communities - if not all, as evident upon looking at some of the modules - is positive, a result implying that the series describing the activity of these areas are concordant for most of the time. For what concerns the links between communities, instead, the picture becomes much more interesting: the third and fourth community in fig. 10 are, in fact, connected by mostly negative links, an evidence suggesting that the two blocks of areas are ‘in counterphase’ for most of the time; while this holds true for the first and third community as well, the first and fourth community are connected by mostly positive links, i.e. even if these blocks of areas are separated, they still seem to work ‘in phase’.
Validated projections. Filtering our projections with a homogeneous benchmark makes the picture even cleaner, as each block is, now, constituted by links of practically the same sign: more precisely, while positive links constitute the majority of links, negative links are found only between modules. Filtering our projections with a heterogeneous benchmark, instead, returns a picture which is halfway between the naïve one and the bSRGM-induced one, as the blocks are sparser than the naïve ones but less homogeneous than the ones induced by the bSRGM. The overall message conveyed by such an analysis, however, does not change: the brain seems to be constituted by groups of areas working either ‘in phase’ or ‘in counterphase’ most of the time, each one being constituted by areas working ‘in phase’ most of the time.
Figure 12: Heatmaps of the communities partitioning the bSCM-induced projections of four subjects. The communities partitioning the bSCM-induced projections of subjects 106016106016106016, 857263857263857263, 115320115320115320 and 654754654754654754 are detected by minimising the Bayesian Information Criterion (BIC). Matrices are coloured according to eq. 96 and eq. 97, i.e. blocks are blue (red) if the majority of links populating them is positive (negative); white blocks on the diagonal indicate singletons - specifically, a cortical region for subject 106016106016106016 and a subcortical region for subject 654754654754654754. All subjects satisfy the statistical variant of the Relaxed Balance Theory, being characterised by blocks with a majority of positive links both on and off the main diagonal - subjects 115320115320115320 and 654754654754654754 are even characterised by one, diagonal block (i.e. a community) populated by a majority of negative links.
@@ VIII.3.2 Traditionally or relaxedly balanced?
Motivated by the last observation, we now employ our benchmarks to probe the patterns of structural (im)balance at the mesoscopic level. We also explicitly notice that speaking of balance theory for brain networks may appear a bit improper: still, this is the only existing framework to interpret our results.
Traditional theory of balance. According to the strong variant of the traditional theory of balance (TBT), the optimal partition of the set of nodes of a network consists of two groups, whose internal (external) connections are dominated by positive (negative) signs; by contrast, the weak variant of the TBT allows for any number of groups. More formally, both versions of the TBT can be probed by minimising the frustration, defined as
@@
i.e. as the number of negative links within modules (indicated with a filled dot) plus the number of positive links between modules (indicated with an empty dot).
Figure 13: Basic statistics of the partitions across subjects. Top panels: histograms of the number of communities detected on the projections returned by each benchmark and optimisation criterion. While $F$ avoids creating frustrated patterns, by individuating a smaller number of larger communities, BIC singles out a larger number of smaller communities. Middle panels: histograms of the size of communities detected on the projections returned by each benchmark and optimisation criterion. These panels refine the picture provided by the previous ones, clarifying that $F$ tends to separate few singletons from the set of remaining nodes. BIC tends to individuate singletons as well, although the size of communities into which the set of remaining nodes is partitioned varies over a broader range (approximately, from 111 to 404040). Bottom panels: number of times each brain region is individuated as a singleton. As evident from the plot, this is mainly observed for subcortical regions, acting as ‘integrators’ of cortical ones. More quantitatively, the fraction of singletons that are identifiable as subcortical regions amounts at i) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.9$\simeq$ 0.9≃ 0.9 according to BIC, on the naïve projections; ii) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.7$\simeq$ 0.7≃ 0.7 according to BIC, on the bSRGM-induced projections; iii) 0.8$\simeq$ 0.8≃ 0.8 according to $F$ and 0.5$\simeq$ 0.5≃ 0.5 according to BIC, on the bSCM-induced projections.
As figs. 10 and 11 show, minimising $F$ always leads to a smaller number of communities with respect to the one revealed by minimising BIC: in fact, from 222 to 444, depending on the projection. The naïve and bSCM-induced ones are the most similar, as 222 mostly positive blocks (one being much larger than the others), separated by a vast majority of negative links, are individuated. The naïve, $F$-induced partition also illustrates the way the optimisation of $F$ works: the large number of positive links leads $F$ to collect them together, being more convenient to accommodate a small number of negative links within a larger, mostly positive, block than letting a large number of positive links be placed between blocks. In any case, the minimisation of $F$ further confirms that none of our projections obeys the traditional formulation of the balance theory.
Statistical theory of balance. Such a result motivates us to inspect the theory of balance best accommodating our observations. To answer such a question, let us, first, notice that the TBT can be formalised upon writing
@@
with $p_{rr}^{+}$ indicating the probability that any two nodes belonging to the same block $r$ are connected by a positive link, $p_{rs}^{+}$ indicating the probability that any two nodes belonging to the different blocks $r$ and $s$ are connected by a positive link and analogously for their negative counterparts. As noticed in Gallo et al. (2024b), however, the positions defining the TBT can be replaced with the milder ones reading
@@
which amounts at requiring $p_{rr}^{+}>p_{rr}^{-}$, $\forall\>r$ and
@@
which amounts at requiring $p_{rs}^{+}<p_{rs}^{-}$, $\forall\>r<s$: in words, the deterministic rules firstly defined by Cartwright, Harary and Davis are replaced by a set of probabilistic criteria individuating ‘a tendency’ to obey, or not to obey, the TBT. More formally, a configuration satisfying eqs. 96 and 97 will be claimed to support the statistical variant of the TBT - specifically, its strong variant if $k=2$ and its weak variant if $k>2$; if, instead, $p_{rr}^{+}\leq p_{rr}^{-}$ for some diagonal blocks or $p_{rs}^{+}\geq p_{rs}^{-}$ for some off-diagonal blocks, it will be claimed to support the statistical variant of the relaxed balance theory (RBT).
Let us, again, take a look at figs. 10 and 11: although $p_{rs}^{+}\geq p_{rs}^{-}$ for all diagonal blocks (i.e. with $r=s$), $p_{rs}^{+}\geq p_{rs}^{-}$ for several, off-diagonal blocks (i.e. with $r\neq s$) as well: we are, thus, led to the conclusion that all projections obey the statistical variant of the RBT Gallo et al. (2024b) as a non-negligible number of positive links is found between - and not only within - clusters.
To evaluate the alignment of our projections with the statistical variant of the TBT, let us count the number of subjects obeying eqs. 96 and 97. When considering the BIC-induced partitions, we find that no subject obeys the statistical variant of the TBT, irrespectively from the benchmark employed: while a vast majority of diagonal blocks is, in fact, populated by a number of positive links that is larger than the number of negative ones - although subjects admitting diagonal blocks populated by a majority of negative links exist (see also fig. 12) - also several, off-diagonal blocks are populated by a number of positive links that is larger than the number of negative ones - as the two examples in figs. 10 and 11 clearly show, $p_{rs}^{+}\geq p_{rs}^{-}$ for many of them.
When considering the $F$-induced partitions, instead, 65 naïve projections, 47 bSRGM-induced projections and 69 bSCM-induced projections obey the statistical variant of the TBT.
Figure 14: Distributions of BIC and $F$ values across the partitions induced by their minimisation. Top panels: histograms of the BIC values characterising the partitions determined by minimising it and Yeo’s partitions. As expected, the former are characterised by a slightly smaller value of such a score than Yeo’s partitions. Middle panels: histograms of the $F$ values characterising the partitions determined by minimising it and Yeo’s partitions. In this second case, the former are characterised by a much smaller value of such a score than Yeo’s partitions. Bottom panels: histograms of the $F$ values characterising the partitions determined by minimising BIC and Yeo’s partitions. In this third case, both are characterised by a quite large value of $F$. Taken together, our results indicate that brain networks do not seek to minimise frustration although BIC minimisation seems (to embody a principle that is) capable of capturing a structure quite similar to the one individuated by Yeo’s task-positive regions. Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule.
@@ VIII.4 Towards a representative brain?
Let us, now, move from individual analyses to the aggregated one, the aim being that of understanding the extent to which the previous results can be generalised.
@@ VIII.4.1 Distribution of the number of communities
First, let us plot the distribution of the number of communities per benchmark and optimisation criterion. The results are shown in fig. 13: the number of communities detected by minimising the frustration is practically always smaller than the number of communities detected by minimising BIC - the $F$-induced mode being 222 and the BIC-induced mode varying from 777 to 888; moreover, the $F$-induced distributions are narrower than the BIC-induced ones, a result hinting at a more pronounced inter-subject variability, in the second case.
@@ VIII.4.2 The role of singletons
Figure 13 illustrates the histograms of the number of communities detected on the projections returned by each benchmark and optimisation criterion, revealing that $F$ individuates a smaller number of larger communities while BIC singles out a larger number of smaller communities. The histograms of the size of communities detected on the same projections refine the aforementioned picture, clarifying that $F$ tends to individuate (overall, too) simple structures such as few singletons detached from the rest of the nodes; although BIC individuates singletons as well, the size of communities into which the rest of the nodes is partitioned ranges from 111 to 404040.
Interestingly, singletons often coincide with the so-called subcortical regions, i.e. areas serving as hubs ‘integrating’ the cortical regions, among the other things: more quantitatively, the fraction of subcortical singletons amounts at i) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.9$\simeq$ 0.9≃ 0.9 according to BIC, on the naïve projections; ii) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.7$\simeq$ 0.7≃ 0.7 according to BIC, on the bSRGM-induced projections; iii) 0.8$\simeq$ 0.8≃ 0.8 according to $F$ and 0.5$\simeq$ 0.5≃ 0.5 according to BIC, on the bSCM-induced projections.
@@ VIII.4.3 Distribution of BIC and $F$ values
A related analysis is the one concerning the distributions of the BIC and $F$ values across the partitions determined by minimising each of them. As fig. 14 confirms, the partitions retrieved by minimising BIC are characterised by a smaller value of such a score than Yeo’s partitions: while this is expected, it should be also noticed that these differences are not very large. Besides, both kinds of partitions are characterised by a similar, and quite large, value of $F$. Inspecting the distribution of the values of $F$, instead, reveals that the partitions retrieved by minimising it are much less frustrated than Yeo’s partitions. Taken together, our results suggest that while brain networks do not seek to minimise frustration, BIC minimisation represents a recipe that is capable of capturing a structure de facto very close to the one captured by Yeo’s task-positive regions.
VIII.4.4 The barycentrical partition
Let us, now, ask ourselves if a criterion to condensate the information embodied by each, individual partition can be devised; naturally, the aforementioned results let us opt for solely focusing on the BIC-induced ones.
A simple recipe could be that of averaging the individual projections in a benchmark-specific fashion; still, the presence of a vast majority of positive links and the spatial incoherence of the negative ones would cause most of the latter to be deleted.
Alternatively, one could try to individuate the ‘most representative’ brain by individuating the ‘most representative’ mesoscale structure: finding an answer to this question ultimately amounts at finding a method to compare any two partitions, say $X\equiv\{x_{1}\dots x_{n}\}$ and $Y\equiv\{y_{1}\dots y_{m}\}$.
Rand index (RI). One of the simplest metric is the RI, reading
@@ VIII.4.4 The barycentrical partition
[…]
Figure 15: Brain centroids individuated by our measures of distance between partitions. Top panels: brain centroids individuated by adopting the Rand Index (RI), the Normalised Mutual Information (NMI), the Rajski distance (RD), the Wallace Index (WI), the Jaccard Index (JI) and the Variation of Information (VI). Generally speaking, while RI, NMI and RD individuate projections with a number of communities lying on the far-right side of the spectrum (i.e. from 999 to 111111), WI, JI and VI individuate projections with a number of communities lying on the far-left side of the spectrum (i.e. 444). Such a result may be explained by noticing that indices like WI and JI are solely defined in terms of true positives: as a consequence, identifying a smaller number of larger communities may enhance the detection of such events. Indices like RI, on the other hand, account for true negatives as well, whence the larger number of smaller communities. Bottom panels: different metrics induces a different overlap between our structure-based partition and Yeo’sfunction-based one: while the larger communities individuated by WI, JI and VI tend to merge large portions of different Yeo’s areas, the smaller ones individuated by RI, NMI and RD tend to be populated by (even a large portion of) single Yeo’s areas.
@@
[…]
@@
[…]
Finding the barycentre. Any of the metrics above can be employed to assess how different the partitions characterising the brains of any two subjects $u$ and $u^{\prime}$ are. Once such a matrix of distances, say $\mathbf{D}\equiv\{D_{uu^{\prime}}\}_{u,u^{\prime}=1}^{U}$, with $U=100$, has been obtained, a second method is needed to condensate such a wealth of information. Our proposal is that of calculating the vector $\underline{\mathbf{D}}\equiv\{\underline{D}_{u}\}$, where
@@
from the others, i.e. at the minimum, average one. Such a partition will be refereed to as to the barycentrical partition or brain centroid. Let us explicitly notice that while VI and RD are proper distances over the space of possible partitions of a system, hence inducing the positions $D_{uu^{\prime}}^{\text{VI}}=\text{VI}_{uu^{\prime}}$ and $D_{uu^{\prime}}^{\text{RD}}=\text{RD}_{uu^{\prime}}$, RI and NMI quantify the similarity of any two partitions [^21] In fact, both RI and NMI could be directly maximised., hence inducing the positions $D_{uu^{\prime}}^{\text{RI}}=1-\text{RI}_{uu^{\prime}}$ and $D_{uu^{\prime}}^{\text{NMI}}=1-\text{NMI}_{uu^{\prime}}$.
The observations about the importance of considering filtered projections and the prominent role played by heterogeneous benchmarks motivate us to focus on the partitions induced by the bSCM.
Upon looking at fig. 15, we realise that i) RI individuates a barycentrical projection with 111111 communities; ii) NMI and RD individuate the same projection, in turn very close to the one individuated by RI - the only difference being the number of communities, now reading 999; iii) WI, JI and VI individuate a barycentrical projection with 444, major communities (plus a number of singletons). To sum up, while RI, NMI and RD individuate projections with a larger number of communities, WI, JI and VI individuate projections with a smaller number of communities. The explanation of such a result may lie in the evidence that indices like WI and JI are solely defined in terms of true positives: as a consequence, identifying a smaller number of larger communities may enhance the detection of such events - although negative connections become under-represented; indices like RI, on the other hand, account for true negatives as well, whence the larger number of smaller communities - connected by a vast majority of negative links.
As a last comment, let us notice that RI, NMI and RD retrieve partitions with a larger BIC than those retrieved by WI, JI and VI.
Let us, now, compare the community structure of the different brain centroids with the partition defined by the Yeo’s regions introduced before. As fig. 15 shows, each metrics induces a different overlap between our structure-based partition and Yeo’s function-based one: still, while the larger communities individuated by WI, JI and VI tend to merge (even) large portions of different Yeo’s areas, some of the smaller ones individuated by RI, NMI and RD tend to be populated by (even a large portion of) single Yeo’s areas. Moreover, WI, JI and VI tend to return singletons, represented by practically empty rows.

### v3

@@ VIII.3.1 A survey on individual projections
Let us start by considering individual partitions, stressing that a large link density does not prevent BIC minimisation from detecting statistically significant mesoscopic structures: its sensitivity to both the number and the sign of the links, in fact, makes it capable of partitioning both denser and sparser configurations.
Naïve projections. As fig. 10 shows, BIC minimisation reveals the presence of $6$ communities: the vast majority of links within communities - if not all, as evident upon looking at some of the modules - is positive, a result implying that the series describing the activity of these areas are concordant most of the time. For what concerns the links between communities, instead, the picture becomes much more interesting: the fourth and fifth communities in fig. 10 are, in fact, connected by mostly negative links, an evidence suggesting that the two blocks of areas are ‘in counterphase’ most of the time; while this holds true for the fifth and sixth communities as well, the first and second communities are connected by mostly positive links, i.e. even if these blocks of areas are separated, they still seem to work ‘in phase’.
Validated projections. Filtering our projections with a homogeneous benchmark makes the picture even clearer, as each block is now constituted by links of practically the same sign: more precisely, while positive links constitute the majority of links, (few) negative links are found only between modules. Filtering our projections with a heterogeneous benchmark, instead, returns a picture that is halfway between the naïve one and the bSRGM-induced one, as the blocks are sparser than the naïve ones, less homogeneous than those induced by the bSRGM, but richer in negative connections than both. The overall message conveyed by such an analysis, however, does not change: the brain seems to be constituted by groups of areas working ‘in phase’ most of the time and ‘in counterphase’ with those of other groups; still, BIC minimisation admits areas working ‘in counterphase’ even if belonging to the same group (see, for example, the patterns highlighted in figs. 7 and 8), as well as areas working ‘in phase’ even if belonging to different groups.
@@ VIII.3.2 Traditionally or relaxedly balanced?
Motivated by the last observation, we now employ our benchmarks to probe the patterns of structural (im)balance at the mesoscopic level. We also explicitly acknowledge that discussing balance theory for brain networks may seem somewhat inappropriate; nonetheless, this is the only existing framework to interpret our results.
Traditional theory of balance. According to the strong variant of the Traditional Balance Theory (TBT), the optimal partition of a network set of nodes consists of two groups, whose internal (external) connections are positive (negative); by contrast, the weak variant of the TBT allows for any number of groups. More formally, both versions of the TBT can be probed by finding the partition $\bm{\sigma}$ that minimises the frustration, defined as
@@
i.e. as the number of negative links within modules (indicated with a filled dot) plus the number of positive links between modules (indicated with an empty dot). Let us also remark that minimising $F$ leads to results that are very similar to the ones obtained upon maximising the signed modularity gallo2024assessing.
The naïve, $F$-induced partition in fig. 10 illustrates a first way the optimisation of $F$ works: a large number of positive links leads $F$ to collect them together, being more convenient to accommodate a small number of negative links within a larger, mostly positive, block than letting a large number of positive links be placed between blocks. The heterogeneous, $F$-induced partition in fig. 10, instead, illustrates a second way the optimisation of $F$ works: a large number of negative links leads $F$ to split them, even at the price of creating many singletons. In any case, the minimisation of $F$ further confirms that none of our projections obeys the traditional formulation of the Balance Theory.
Statistical theory of balance. Such a result motivates us to inspect the theory of balance best accommodating our observations. To answer such a question, let us first notice that the TBT can be formalised upon writing
@@
with $p_{rr}^{+}$ indicating the probability that any two nodes belonging to the same block $r$ are connected by a positive link, $p_{rs}^{+}$ indicating the probability that any two nodes belonging to the different blocks $r$ and $s$ are connected by a positive link and analogously for their negative counterparts. As noticed in gallo2024assessing, however, the positions defining the TBT can be replaced with the milder ones reading
@@
which amounts to requiring $p_{rr}^{+}>p_{rr}^{-}$, $\forall\>r$ and
@@
which amounts to requiring $p_{rs}^{+}<p_{rs}^{-}$, $\forall\>r<s$: in words, the deterministic rules firstly defined by Cartwright, Harary and Davis are replaced by a set of probabilistic criteria individuating ‘a tendency’ to obey, or not to obey, the TBT. More formally, a configuration satisfying eqs. 97 and 98 will be claimed to support the statistical variant of the TBT - specifically, its strong variant if $k=2$ and its weak variant if $k>2$; if, instead, $p_{rr}^{+}\leq p_{rr}^{-}$ for some diagonal blocks or $p_{rs}^{+}\geq p_{rs}^{-}$ for some off-diagonal blocks, it will be claimed to support the statistical variant of the Relaxed Balance Theory (RBT).
In order to evaluate the alignment of our projections with either statistical variants, let us take a look at the top panels of fig. 10, depicting the BIC-induced partitions: since $p_{rs}^{+}\leq p_{rs}^{-}$ for several, diagonal blocks (i.e. with $r=s$) and $p_{rs}^{+}\geq p_{rs}^{-}$ for several, off-diagonal blocks (i.e. with $r\neq s$), we are led to the conclusion that all projections obey the statistical variant of the RBT gallo2024assessing as a non-negligible number of negative (positive) links is found within (between) - and not only between (within) - clusters. As fig. 11 further confirms, no subject obeys the statistical variant of the TBT, regardless of the benchmark employed, when considering the BIC-induced partitions.
Let us, now, take a look at the bottom panels of fig. 10, depicting the $F$-induced partitions: although eqs. 97 and 98 are not strictly obeyed, our subject, now, admits blocks with a majority of negative links only off the main diagonal and blocks with a majority of positive links only on the main diagonal; we are, thus, led to the conclusion that all projections align better with the statistical variant of the TBT. As fig. 11 further confirms, all subjects align better with the statistical variant of the TBT, regardless of the benchmark employed, when considering the $F$-induced partitions.
@@ VIII.4 Towards a representative brain?
Let us now move from individual analyses to the aggregated one, the aim being to understand the extent to which the previous results can be generalised.
Figure 12: Basic statistics of the partitions across subjects. Top panels: histograms of the number of communities detected on the projections returned by each benchmark and optimisation criterion. For what concerns naïve and bSRGM-induced projections, $F$ avoids creating frustrated patterns by individuating a smaller number of larger communities, while BIC singles out a larger number of smaller communities; for what concerns bSCM-induced projections, instead, $F$ individuates a larger number of smaller communities than BIC. Middle panels: histograms of the size of communities detected on the projections returned by each benchmark and optimisation criterion. These panels refine the picture provided by the previous ones, clarifying that $F$ tends to separate the singletons from the set of remaining nodes; BIC tends to individuate singletons as well, although the size of communities into which the set of remaining nodes is partitioned varies over a broader range. Bottom panels: number of times each brain region is individuated as a singleton. As evident from the plot, this is almost solely observed for subcortical regions, especially on $F$-induced partitions; the most ‘integrated’ picture is, instead, returned by the bSCM-induced projections, partitioned by minimising BIC.
@@ VIII.4.1 Distribution of the number of communities
First, let us plot the distribution of the number of communities per benchmark and optimisation criterion. The results are shown in fig. 12: for what concerns naïve and bSRGM-induced projections, the number of communities detected by minimising the frustration is practically always smaller than the number of communities detected by minimising BIC - the $F$-induced mode being $2$ and the BIC-induced mode varying from $7$ to $8$; moreover, the $F$-induced distributions are narrower than the BIC-induced ones, a result hinting at a more pronounced inter-subject variability, in the second case. Concerning the bSCM-induced projections, instead, $F$ identifies a larger and more dispersed number of smaller communities than BIC.
@@ VIII.4.2 The role of singletons
The histograms of the size of communities detected on the same projections refine the aforementioned picture, clarifying that $F$ tends to individuate simpler structures than BIC, such as fewer modules accompanied by several singletons detached from the rest of the nodes - although BIC may individuate singletons as well, the size of communities into which the rest of the nodes is partitioned varies over a broader range.
Figure 13: Distributions of the BIC and $F$ values across the partitions induced by their minimisation. Top panels: histograms of the BIC values characterising the partitions determined by minimising it (in blue) and Yeo’s partitions (in orange). As expected, the former are characterised by a slightly smaller value of such a score than Yeo’s partitions. Middle panels: histograms of the $F$ values characterising the partitions determined by minimising it (in blue) and Yeo’s partitions (in orange). In this second case, the former are characterised by a (much) smaller value of such a score than Yeo’s partitions. Bottom panels: histograms of the $F$ values characterising the partitions determined by minimising BIC (in blue) and Yeo’s partitions (in orange). In this third case, both are characterised by a quite large value of $F$. Bin widths are computed according to the Freedman-Diaconis rule. Taken together, our results indicate that i) brain networks do not seek to minimise frustration - even when removing the subcortical areas; ii) BIC minimisation embodies a principle returning a structure that captures (at least) part of the features of the Yeo’s task-based one.
Interestingly, singletons often coincide with the so-called subcortical regions, i.e. areas serving as hubs ‘integrating’ the cortical regions, among the other things: more quantitatively, the fraction of subcortical singletons amounts to i) $\simeq 0.86$ according to $F$ and $\simeq 0.79$ according to BIC, on the naïve projections; ii) $\simeq 0.87$ according to $F$ and $\simeq 0.59$ according to BIC, on the bSRGM-induced projections; iii) $\simeq 0.74$ according to $F$ and $\simeq 0.03$ according to BIC, on the bSCM-induced projections. In words, carrying out a BIC-based community detection on bSCM-induced projections rarely leads to isolate subcortical areas, a result indicating that accounting for local properties such as the series- and the time-specific degrees leads to recover projections where all regions are (practically always) integrated.
@@ VIII.4.3 Distribution of BIC and $F$ values
A related analysis concerns the distributions of the BIC and $F$ values across the partitions determined by minimising each of them. As fig. 13 confirms, the partitions retrieved by minimising BIC are characterised by a smaller value of such a score than Yeo’s partitions; besides, both kinds of partitions are characterised by a similar, and quite large, value of $F$. Inspecting the distribution of the values of $F$, instead, reveals that the partitions retrieved by minimising it are much less frustrated than Yeo’s partitions. Taken together, our results suggest that i) brain networks do not seek to minimise frustration (even when removing subcortical areas - as in this case, to carry out a fair comparison with Yeo’s classification); ii) BIC minimisation is capable of capturing a structure that shares several similarities with the one defined by Yeo’s task-based regions.
Figure 14: Brain centroids individuated by the considered distances between partitions. Top panels: brain centroids individuated by adopting the Rand Index (RI), the Normalised Mutual Information (NMI), the Rajski distance (RD), the Wallace Index (WI), the Jaccard Index (JI), and the Variation of Information (VI). Generally speaking, while RI, NMI, and RD individuate projections with a number of predominantly positive communities lying on the right side of the spectrum (i.e. from $7$ to $10$), WI, JI, and VI individuate projections with a number of predominantly positive communities lying on the left side of the spectrum (e.g. 55). Such a result may be explained by noticing that indices like WI and JI are solely defined in terms of true positives: as a consequence, identifying a smaller number of larger communities may enhance the detection of such events. Indices like RI, on the other hand, account for true negatives as well, whence the larger number of smaller communities. Bottom panels: different metrics induce a different overlap between our structure-based partitions and Yeo’s task-based one. Overall, the communities identified by WI, JI, VI, and RD are able to capture up to $75\%$ of certain Yeo’s regions, while the smaller ones identified by RI and NMI rarely capture more than $50\%$ of a given Yeo’s region.
VIII.4.4 The barycentric partition
Let us, now, ask ourselves if a criterion to condense the information embodied by each individual partition can be devised. To this aim, one could try to single out the ‘most representative’ brain by individuating the ‘most representative’ mesoscale structure: finding an answer to this question ultimately amounts to finding a method to compare any two partitions, say $X\equiv\{x_{1}\dots x_{n}\}$ and $Y\equiv\{y_{1}\dots y_{m}\}$.
Rand index (RI). One of the simplest metrics is the RI, reading
@@ VIII.4.4 The barycentric partition
and computing the ratio between the sum of true positives (TP; the number of nodes that are ‘together’ in both partitions, i.e. in the same subset in $X$ and in the same subset in $Y$) and true negatives (TN; the number of nodes that are ‘separated’ in both partitions, i.e. in different subsets in $X$ and in different subsets in $Y$) and the total number of pairs of nodes. Naturally, $0\leq\text{RI}\leq 1$.
Other variants are the Wallace Index (WI), defined as $\text{WI}=\text{TP}/(\text{TP}+\text{FP})$, and the Jaccard Index (JI), defined as $\text{JI}=\text{TP}/(\text{TP}+\text{FP}+\text{FN})$: notice, however, that both WI and JI measure the agreement between two partitions by solely considering the number of TP.
Figure 15: Average brain, partitioned according to BIC minimisation and $F$ minimisation. As the presence of a vast majority of positive links causes most of the negative ones populating the naïve and bSRGM-induced projections to be deleted, only the bSCM-induced projections have been considered: while the average brain partitioned according to $F$ minimisation is characterised by $2$ predominantly positive communities that gather the vast majority of Yeo’s regions together and 66 predominantly negative modules that fragment the remaining ones, the average brain partitioned according to BIC minimisation is characterised by modules sharing similarities with the ones induced by NMI and those induced by VI, and capable to capture up to the $80\%$ of certain Yeo’s regions.
@@
with $h_{ij}=|X_{i}\cap Y_{j}|/N$ being the fraction of nodes common to subsets $X_{i}$ and $Y_{j}$, $S(X)=-\sum_{i=1}^{N}f_{i}\ln f_{i}$ being the Shannon entropy induced by the partition $X$ and $f_{i}=|X_{i}|/N$ being the fraction of nodes within the cluster $X_{i}$, $S(Y)=-\sum_{j=1}^{N}g_{j}\ln g_{j}$ being the Shannon entropy induced by the partition $Y$ and $g_{j}=|Y_{j}|/N$ being the fraction of nodes within the cluster $Y_{j}$. Naturally, $0\leq\text{NMI}\leq 1$.
@@
naturally, $0\leq\text{RD}\leq 1$.
Finding the barycentre. Any of the metrics above can be employed to assess how different the partitions characterising the brains of any two subjects, say $u$ and $u^{\prime}$, are. Once such a matrix of distances, say $\mathbf{D}\equiv\{D_{uu^{\prime}}\}_{u,u^{\prime}=1}^{U}$, with $U=100$, has been obtained, a second method is needed to condense such a wealth of information. Our proposal is that of calculating the vector $\underline{\mathbf{D}}\equiv\{\underline{D}_{u}\}$, where
@@
from the others, i.e. at the minimum, average one. Such a partition will be referred to as the barycentric partition or brain centroid. Let us explicitly notice that while VI and RD are proper distances over the space of possible partitions of a system, hence inducing the positions $D_{uu^{\prime}}^{\text{VI}}=\text{VI}_{uu^{\prime}}$ and $D_{uu^{\prime}}^{\text{RD}}=\text{RD}_{uu^{\prime}}$, RI and NMI quantify the similarity of any two partitions [^20] In fact, both RI and NMI could be directly maximised., hence inducing the positions $D_{uu^{\prime}}^{\text{RI}}=1-\text{RI}_{uu^{\prime}}$ and $D_{uu^{\prime}}^{\text{NMI}}=1-\text{NMI}_{uu^{\prime}}$.
The observations about the importance of considering filtered projections and the prominent role played by heterogeneous benchmarks motivate us to focus on the partitions obtained by minimising BIC on the projections induced by the bSCM.
Upon looking at fig. 14, we realise that i) RI individuates a barycentric projection with $10$ predominantly positive communities; ii) NMI individuates a barycentric projection with $8$ predominantly positive communities; iii) RD individuates a barycentric projection with $7$ predominantly positive communities; iv) WI, JI, and VI individuate a barycentric projection with $5$, major communities defined by predominantly positive connections. The explanation of such a result may lie in the evidence that indices like WI and JI are solely defined in terms of true positives: as a consequence, identifying a smaller number of larger communities may enhance the detection of such events - although negative connections become under-represented; indices like RI, on the other hand, account for true negatives as well, whence the larger number of smaller communities - connected by a vast majority of negative links.
Let us, now, compare the community structure of the different brain centroids with the partition defined by Yeo. As fig. 14 shows, each metrics induces a different overlap between our structure-based partitions and Yeo’s task-based one: while the communities individuated by WI, JI, VI, and RD tend to capture (even) large portions of certain Yeo’s communities - up to the $75\%$ for the first three indices and up to the $60\%$ for the fourth one - the smaller ones individuated by RI and NMI rarely capture more than half of a given Yeo’s community.
A simpler recipe to identify the representative brain could be that of averaging the individual projections in a benchmark-specific fashion; as the presence of a vast majority of positive links causes most of the negative ones populating the naïve and bSRGM-induced projections to be deleted, here we limit ourselves to consider the bSCM-induced projections. As fig. 15 shows, both the BIC minimisation and the $F$ minimisation lead to single out $8$ communities: in the second case, these modules appear as quite uninformative as $6$ are very small groups of nodes fragmenting a single Yeo’s region in a quite unrealistic fashion, while the seventh and the eighth ones gather all the remaining regions together; in the first case, instead, these modules share similarities with the ones induced by NMI and those induced by VI, and are capable to capture up to the $80\%$ of certain Yeo’s regions.